OpenStudy (anonymous):

andrew can paint the neighbor's house 4 times as fast as bailey. The year Andrew and bailey worked together, it took 9 days. How long would it take each to paint the house? I need help on this Please

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

set it up as an equation.

OpenStudy (anonymous):

andrew can paint two times as fast as bailey so to represent andrew you can put 2B. as an equation you can put A+B=9 you know A=2B 2B+B=9 3B=9 B=3 so bailey is 3 and then plug back in B to a=2B A=2(3) A=6 hope this helps!